OpenStudy (anonymous):

How many 4-letter permutaions can you form from the letters of the word, "MODEL"?

OpenStudy (anonymous):

Do they have to be valid words? Otherwise (5!)/(1!)

OpenStudy (anonymous):

you can't use one letter more than once for first letter , you have 5 choices second , 4 third , 3 second , 2 first , 1 multiply them all 5*4*3*2*1=

OpenStudy (anonymous):

It's four-lettered words so it would just be 5*4*3*2 wouldn't it? Even though it's the same answer the fifth (and last) letter does technically not get used.

OpenStudy (anonymous):

makes sense thank you!

OpenStudy (anonymous):

what if the word is LINEAR so it has 6 letters but you need to make 4 letter words, how would you do that?

OpenStudy (anonymous):

Then it's 6!/(6-4)!=6!/2! It's n!/p! where n is the number of choices available (in this case 6) and p is the number of choices available minus the number of open spots (in this case 6-4=2)